HUBBLE SPACE TELESCOPE DAILY REPORT #3086 PERIOD COVERED: 0000Z (UTC) 04/02/02 - 0000Z (UTC) 04/03/02 Daily Status Report as of 093/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.2 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.3 Completed Six Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.4 Completed Two Sets of FGS/1 9234 (Calibrating the Mass-Luminosity Relation at the End of the Main Sequence) Fine Guidance Sensor #1R was used to calibrate the mass-luminosity relation for stars less massive than 0.2 Msun, with special emphasis on objects near the stellar/brown dwarf border. The proposal completed with no reported errors. 1.5 Completed Two Sets of ACS/WFC/HRC 8992 (ACS Early Release Observations) The Advanced Camera for Surveys (WFC and HRC) was used to perform early release observations {ERO}, employing a variety of instruments to demonstrate the new capabilities. The observations (not all in this iteration) will include: 1} WFC observations of peculiar galaxies NGC 4676 {The Mice} and UGC 10214. This will be used to study young stellar clusters formed through the process of galaxy interaction. Any dark companion to UGC 10214 will also be searched for in sensitive ACS/WFC observations. 2} Emission line and continuum imaging of the Cone Nebula and M 17 with WFC to demonstrate the improved sensitivity and angular resolution of ACS. 3} Coronagraphic imaging of HD 141569, a main sequence star with circumstellar disk. 4} Coronagraphic imaging with the HRC of the nearby quasar 3C 273 to investigate the physical properties of the host galaxy and environments. There were no reported problems. 1.6 Completed WF/PC-2 8952 (SM3B Flat Field Calibration) The WF/PC-2 was used to check for changes in flat field using external earth flat fields. No problems were encountered. 1.7 Completed Three Sets of FGS/2/3 8995 (Verification of Operational FGS Calibrations) Fine Guidance Sensors #2 and #3 were used to monitor the characteristics of HST's Fine Guidance Sensors across the boundary of the servicing mission SM3B and compares the data with pre-SM3B baseline characterizations. For each FGS, stars from a selected field in the standard astrometric Galactic cluster M35 are observed in Position mode. This provides a measurement of their relative positions, which in turn provides an effective means to detect and monitor any changes in the FGS optical distortions and plate scales. These observations will obtain S-curves at the center of each FGS FOV. If changes larger than tolerance are observed, it will be necessary to activate the appropriate contingency proposal that will obtain data for new S-curves. These data will also be used to re-compute the optimal values of parameters, or K-factors, that are used in the acquisition and tracking of guide stars {and astrometry targets}. The proposal completed with no reported problems. 1.8 Completed WF/PC-2 9258 (Observatory Focus Monitor) The WF/PC-2 was used to more accurately define the HST focus. The HST focus drifts slowly and shows evidence of undergoing slips of a few microns at random times. The rate of the WF/PC-2 monitoring program is insufficient to track and/or understand OTA behavior in order to request timely and appropriate Secondary Mirror corrections. There were no reported problems. 1.9 Completed STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 12 Successful: 12 Scheduled Re-acquisitions: 4 Successful: 4 2.2 FHST Updates: Scheduled: 19 Successful: 19 2.3 Operations Notes: The NICMOS Cryo Cooler cool down is nominally proceeding. Using ROP SR-9A, the SSR-3 EDAC error counter was cleared at 092/1403Z. The Fine Corrector Alignment, Part 2 was completed at 092/1604Z as directed by an operations request. Using the data from the second visit of the ACS Fine Alignment proposal (9014), the best corrector settings were determined. The corrector was moved to the new positions so that as many science exposures as possible will be taken at these best settings. SSA transmitter #2 was turned on for the interval 092/2141Z to 092/2210Z per ROP IC-2. A re-planned intercept SMS in the process of being prepared. This SMS will modify two ACS Early Release Observation proposals since previous iterations had unusable faint guide stars. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW